OpenStudy (anonymous):

What is the general form of the equation for the given circle? x2 + y2 − 8x − 8y + 23 = 0 x2 + y2 − 8x − 8y + 32 = 0 x2 + y2 − 4x − 4y + 23 = 0 x2 + y2 + 4x + 4y + 9 = 0

OpenStudy (anonymous):

idk how to put a pic on here so im just gonna describe the circle

OpenStudy (anonymous):

it's 90 degrees with points A (4,7) and B (7,4)

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

@Marcorie

OpenStudy (anonymous):

A) (x-4)^2 + (y-4)^2 = 3

OpenStudy (anonymous):

do i solve that or something?

OpenStudy (anonymous):

No i solved the first equation this is standard form for a circle

OpenStudy (anonymous):

(x-h)^2 + (y-h)^2 = r^2

OpenStudy (xapproachesinfinity):

you have to show how ? no direct answer are allowed in OP

OpenStudy (anonymous):

Okay sorry I'm new to this

OpenStudy (anonymous):

i dont have to show how i just need the answer @xapproachesinfinity

OpenStudy (anonymous):

(x-h)^2 + (y-h)^2 = r^2

OpenStudy (anonymous):

it's fine i just really need the answer @Deezzz

OpenStudy (xapproachesinfinity):

it is not allowed in open study!!

OpenStudy (anonymous):

where (h,k) is the center and r is the radius of the circle

OpenStudy (anonymous):

please just help!! i honestly dont care about the work i just really need the answer so i could pass this damn test

OpenStudy (xapproachesinfinity):

whether you just need the answer or not you have the follow the guide lines

OpenStudy (anonymous):

-.- @xapproachesinfinity

OpenStudy (anonymous):